Can anyone help me solve this math problem? See below.
An urn contains 2-one dollar bills, 1 five dollar bill and 1 ten dollar bill. A player draws bills one at a time without replacement from the urn until a ten-dollar bill is drawn. Then the game stops. All bills are kept by the player. Determine:
(A) The probability of winning $11
(B) The probability of winning all bills in the urn.
(c) The probability of the game stoping at the second draw.

Answer by stanbon(75887)

Can anyone help me solve this math problem? See below.
An urn contains
2-one dollar bills,
1 five dollar bill and
1 ten dollar bill.
---
A player draws bills one at a time without replacement from the urn until a ten-dollar bill is drawn. Then the game stops. All bills are kept by the player. Determine:
(A) The probability of winning $11
How?: $1 then $10
Probability: (2/5)(1/4) = 1/10
-----------------------------------
(B) The probability of winning all bills in the urn.
How?: not 10, then not 10, then not ten, then 10
Probability: (1-(1/4))(1-(1/3))(1-(1/2)1
= (3/4)(2/3)(1/2)
= 1/4
------------------------------------
(c) The probability of the game stopping at the second draw.
How? anything but 10 followed by 10
Probability: (3/4)(1/3) = 1/4
===============================
Cheers,
Stan H.